If 180 is divided by x, then the remainder is 2. What is the remainder when 357 is divided by x
Answer by htmentor(1343)

Let n = the whole number when 180 is divided by x.
Then we can write 180/x = n + 2/x since the remainder is 2
Since 357 = 2*180 - 3, 357 divided by x is:
(2*180 - 3)/x = 2*180/x - 3/x
But 180/x = n + 2/x -> 2(n+2/x) - 3/x = 2n + 4/x - 3/x = 2n + 1/x
2n is a whole number and we are left with the remainder of 1.
Ans: 1
Take an example:
180/89 = 2 + 2/89 [remainder 2]
357/89 = 4 + 1/89 [remainder 1]